# Distillation of Large Language Models via Concrete Score Matching

### v1

@@ Abstract
Large language models (LLMs) deliver remarkable performance but are costly to deploy, motivating knowledge distillation (KD) for efficient inference. Existing KD objectives typically match student and teacher probabilities via softmax, which blurs valuable logit information. While direct logit distillation (DLD) mitigates softmax smoothing, it fails to account for logit shift invariance, thereby restricting the solution space. We propose Concrete Score Distillation (CSD), a discrete score-matching objective that overcomes both softmax-induced smoothing and restrictions on the optimal solution set. We resolve the training instability and quadratic complexity of discrete score-matching in autoregressive LLMs, and the resulting CSD objective aligns relative logit differences across all vocabulary pairs between student and teacher with flexible weighting. We provide both mode-seeking and mode-covering instances within our framework and evaluate CSD on task-agnostic instruction-following and task-specific distillation using GPT-2-1.5B, OpenLLaMA-7B, and GEMMA-7B-IT. Experiments show that CSD consistently surpasses recent KD objectives, achieves favorable fidelity–diversity trade-offs, and yields complementary gains when combined with on-policy techniques, demonstrating its scalability and effectiveness for LLM distillation.
@@ 1 Introduction
In traditional KD, direct logit distillation (DLD) (Ba & Caruana, 2014; Urban et al., 2017) has been proposed as an alternative strategy, with advantages in generalization capability and in removing the softmax smoothing (Kim et al., 2021). However, such approaches have not been thoroughly explored in the context of LLMs. This paper identifies a key drawback of DLD: its restriction on the optimal solution set as described in Figure˜1(c). Considering the softmax activation in inference, it is sufficient for the teacher’s and student’s logits to agree up to an additive constant, but the previous solutions of DLD fail to accommodate such an acceptable slack constant, a.k.a. logit shift invariance. Such a restriction on the solution set may hinder the discovery of optimal solutions in distillation, particularly when the teacher and student models have a large capacity gap, as is often the case with LLMs. Therefore, the goal of this paper is to establish a design space of distillation losses that overcome both the softmax-induced smoothing of teacher knowledge and the restriction on the solution set.
This paper adopts the idea from energy-based models (Song & Kingma, 2021), which design objectives that avoid the constraint of probabilistic models (sum-to-one) by using the score-matching objective (Hyvärinen & Dayan, 2005). We propose Concrete Score Distillation (CSD), a discrete form of the score-matching objective (Meng et al., 2022) adapted for autoregressive LLM distillation. We address training instability and computational overhead arising when applying the score-matching objective to LLMs, and provide theoretical guarantees of optimality, showing that its solution set is broader than that of DLD from both theoretical and empirical perspectives. The resulting objective reduces to matching the relative logit differences across all pairs of vocabulary items between the student and teacher, while allowing flexible weighting across all vocabulary pairs in linear time with respect to vocabulary size. Furthermore, we present instances within our framework that exhibit both mode-seeking and mode-covering properties.
In our experiments, we conducted both task-agnostic instruction-following distillation and task-specific distillations (summarization, mathematics, and translation) using GPT-2 (Radford et al., 2019), OpenLLaMA (Geng & Liu, 2023), and GEMMA (Team et al., 2024) backbones. The proposed CSD consistently outperformed recent probability-matching objectives as well as direct logit distillation. By appropriately choosing weighting functions, we further demonstrated that our method resides on the frontier of the diversity–fidelity trade-off. Finally, we observed complementary performance gains when integrating our loss with on-policy techniques.
@@ 2.1 Knowledge Distillation of Large Language models
We consider autoregressive large language models (LLMs), consisting of a teacher $p_{T}$ and a student $q_{\theta}$, where the student is a smaller and more efficient model. Given an input context $\mathbf{c}$}, the student generates an output sequence ${\mathbf{y}}=(y_{1},y_{2},\dots,y_{L})$ with probability $q_{\theta}({\mathbf{y}}|{\mathbf{c}})=\prod_{t=1}^{L}{q_{\theta}(y_{t}|{\mathbf{c}},{\mathbf{y}}_{<t})}$, where $L$ denotes the sequence length, and the teacher’s probability is defined analogously. Each token $y_{t}$ is drawn from the fixed vocabulary set $\mathcal{V}:=\{v_{1},v_{2},$...\}. As in prior works (Lin et al., 2020; Ko et al., 2024), we assume the teacher and student share the same vocabulary set. To compute the token probability $q_{\theta}(y_{t}|{\mathbf{c}},{\mathbf{y}}_{<t})$, an LLM typically adopts a parametric function $f_{\theta}:{\mathcal{V}}^{|{\mathbf{c}}|}\times{\mathcal{V}}^{t-1}\rightarrow\mathbb{R}^{|{\mathcal{V}}|}$, which maps the input $({\mathbf{c}},{\mathbf{y}}_{<t})$ to a logit vector $f_{\theta}({\mathbf{c}},{\mathbf{y}}_{<t})\in\mathbb{R}^{|{\mathcal{V}}|}$. The logit corresponding to token $y_{t}$ is denoted by $f_{\theta}({\mathbf{c}},{\mathbf{y}}_{<t})[y_{t}]$. For brevity of notation, the explicit input arguments of the function $f_{\theta}$ will be omitted hereafter. Let $f_{T}$ be the parametric function of the teacher. Accordingly, the probability of each token is calculated through the following softmax transformation:
@@
However, $D_{\text{KL}}$ focuses on the teacher’s probabilities and is constrained by the softmax. As shown in Figure˜1(b), although the teacher carries rich knowledge across all vocabulary items at the logit level, much of it is lost after softmax, and the teacher provides nearly identical gradient signals to most minor tokens. Accordingly, in classical KD studies (Ba & Caruana, 2014; Urban et al., 2017), direct logit distillation (DLD) has been widely adopted as a logit-level mean squared error (MSE) loss:
@@ 2.2 Score Matching for a discrete random variable
Score-matching (SM) (Hyvärinen & Dayan, 2005) was originally proposed in energy-based models (Song & Kingma, 2021) with continuous variables ${\mathbf{x}}\in\mathbb{R}^{d}$. An energy function $E_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ maps $\mathbf{x}$} to a scalar. The corresponding probability and the score-matching objective are given by:
@@
where $Z_{\theta}=\int_{{\mathbf{x}}}\text{exp}(-E_{\theta}({\mathbf{x}}))\text{d}{\mathbf{x}}$ is the partition function, and $w(\cdot)$ is a weighting function. The term $\nabla_{{\mathbf{x}}}\log{q_{\theta}({\mathbf{x}})}=-\nabla_{{\mathbf{x}}}E_{\theta}({\mathbf{x}})$ is known as the Stein score, which uniquely identifies the probability distribution without requiring the computation of $Z_{\theta}$. $\mathcal{L}_{\text{SM}}$ facilitates the design of losses without considering the normalization constraint of probabilistic models. The probability computation $q_{\theta}$ here follows, analogously, the form of the LLM probabilities in Eq.˜1. The difference is that an LLM outputs energy values $f_{\theta}$ over all finite states at once, whereas an EBM handles continuous variables, so that each input to $E_{\theta}$ yields only a single scalar output.
@@
where $w(\cdot,\cdot)$ is a positive weighting function. Previous work on language models (Lou et al., 2024) typically adopted this loss by directly parameterizing the concrete score (also known as discrete diffusion models) to mimic the data distribution. In contrast, we take this concept as a starting point to design logit-level distillation losses for autoregressive-type language models, which are more dominant in real-world applications.
@@ 3.1 Concrete Score Distillation for Large Language Models
Training instability is a well-known issue in likelihood ratio estimation (Rhodes et al., 2020). Following Higuchi & Suzuki (2025), we address it by applying a monotonically increasing function to the concrete scores. In particular, we adopt the logarithm, which yields the following objective:
@@
(Consistency) Given context $\mathbf{c}$} and prefix ${\mathbf{y}}_{<t}$, assume sufficient model capacity. For any $w(\cdot,\cdot)>0$, define the set of optimal parameters as $\Theta_{\text{CSD}}^{*}=\operatorname*{arg\,min}_{\theta}{\mathcal{L}_{\text{CSD}}\left(\theta;p_{T},w\right)}$. Then, for any $\theta^{*}\in\Theta_{\text{CSD}}^{*}$, we have $\mathcal{L}_{\text{CSD}}\left(\theta^{*};p_{T},w\right)=0$, and the following holds for all $y_{t}\in\mathcal{V}$:
@@
(Solution Superset) Under sufficient model capacity, let the set of optimal parameters $\Theta_{\text{CSD}}^{*}=\operatorname*{arg\,min}_{\theta}{\mathcal{L}_{\text{CSD}}\left(\theta;p_{T},w\right)}$ and $\Theta_{\text{DLD}}^{*}=\operatorname*{arg\,min}_{\theta}{\mathcal{L}_{\text{DLD}}\left(\theta;p_{T},w\right)}$, then following holds:
@@ 3.2 Gradient Computation and Analysis
1 Compute the student logit $f_{\theta}[y_{t}]=f_{\theta}({\mathbf{c}},{\mathbf{y}}_{<t}),\forall y_{t}\in\mathcal{V}$.
@@
3  Compute the teacher logit $f_{T}[y_{t}]=f_{T}({\mathbf{c}},{\mathbf{y}}_{<t}),\forall y_{t}\in\mathcal{V}$.
@@ 7   Compute the weighted normalized logits:
The proof is provided in Section˜A.3. These results follow from factorizing the independent variables. Algorithm˜1 further details the gradient computation of Eq.˜9 step by step, with each step requiring only linear time over the vocabulary. An alternative approach is to use Monte Carlo estimation. Instead of taking a weighted sum over all possible states of $y_{t}$ with $w_{1}$, one can draw a single sample of $y_{t}$ according to probability $w_{1}$ and compute the loss in expectation. However, this approach increases the variance within the batched samples, leading to degraded estimation accuracy and typically lower performance compared to the analytic computation (see Figure˜5).
@@ 4 Experiments
This section comprehensively validates the effectiveness of the proposed Concrete Score Distillation (CSD) across various experimental setups. Section˜4.1 shows results on task-agnostic instruction-following distillation, comparing CSD with alternative loss functions and assessing its performance when combined with recent on-policy methods. Section˜4.2 further examines task-specific settings, including math, summarization, and translation, to evaluate the applicability of CSD. Finally, Section˜4.3 establishes the contribution of each component in CSD through ablation studies.
@@ 4.1 Task-Agnostic Instruction-Following Distillation
Experimental setup: We follow the training setup of DistiLLM (Ko et al., 2024). For the distillation dataset $\mathcal{D}$, we use databricks-dolly-15k (Conover et al., 2023), containing about 14,000 samples for training, with 500 held out for validation and 500 for evaluation. For comparison with the baseline, we optionally add a pretraining loss using the pretraining dataset OpenWebText (Gokaslan & Cohen, 2019) in some cases of Table˜2. We first fine-tune the GPT-2-1.5B (Radford et al., 2019) teacher on the dataset, and then distill it into GPT-2-0.1B and GPT-2-0.3B students. Similarly, we distill OpenLLaMA-7B (Geng & Liu, 2023) into OpenLLaMA-3B. We determined the learning rate and batch size by referring to the search ranges used in prior studies (Gu et al., 2024; Ko et al., 2024). We use the detached student probability as the default choice for both $w_{1}$ and $w_{2}$, and analyze alternative choices through ablation studies. Please refer to Appendix˜C for further details.
@@
(b) Temperature adjustment in the range $[0.2,1.8]$.
@@
Loss-level comparison: To purely analyze the effect of the distillation loss itself, this comparison excludes the use of pretraining losses, initialization with an SFT-tuned student, and any on-policy techniques. Table˜1 shows that the proposed CSD objective outperforms the other nine objectives, ranking first on three of the five benchmarks, second on one, and achieving the highest average score. SKL (Ko et al., 2024) and AB (Wang et al., 2025) exhibit slightly lower performance than previously reported, likely due to their reliance on pretraining losses or on-policy techniques.
Figure˜3(a) shows the fidelity–diversity trade-off based on ROUGE-L and Self-BLEU scores. Traditionally, KL favors diversity, whereas RKL favors mode-seeking. Within this trade-off, SKL, SRKL, TV, and AB achieve higher ROUGE-L scores than RKL, but at the cost of reduced diversity, reflecting a stronger emphasis on fidelity. Diversity, however, remains an important aspect of user experience in instruction-following, and it becomes a valuable metric as it enhances performance when combined with best-of-N sampling. The proposed CSD provides an additional lever to control the fidelity–diversity trade-off. By default, using the detached student probabilities ($S,S$) yields the highest fidelity. Replacing one side with uniform ($U,S$) or with the teacher ($T,S$) gradually increases diversity. This is likely because the ($S,S$) makes the model focus only on regions where the student already assigns a high likelihood, limiting its exploratory ability. The trade-off offered by CSD envelopes those of existing losses, and we expect that even better operating points may exist within the design space of $w_{1}$ and $w_{2}$. Figure˜3(b) presents an ablation on temperature, which enables easy adjustment of the trade-off during inference. Even within a reasonable range of decoding temperature, CSD achieves better trade-off points than other losses. In particular, CSD ($U,S$) demonstrates a well-balanced exchange between diversity and fidelity through temperature adjustment.
Table 2: Instruction-following performance of CSD with on-policy techniques for various backbones. $\mathcal{D}$ denotes the distillation dataset. ROUGE-L scores are averaged over five random seeds, with the best score for each student highlighted in bold.
@@
Table 3: Task-specific distillation performance from the GEMMA-7B-IT teacher to the GEMMA-2B-IT student.
Summarization Translation GSM8K Loss ROUGE-L COMET Accuracy Teacher 37.09 79.23 60.27 KL 35.02 73.96 24.03 JS 35.60 74.05 23.73 TV 27.49 73.73 0.00 Jeffrey 35.29 74.02 23.28 SKL 25.86 59.65 0.00 SRKL 26.68 73.10 0.00 CSD (Ours) 35.67 74.14 25.78
Orthogonal improvement with recent on-policy advances: Table˜2 reports the performance of recent distillation baselines augmented with the CSD loss, demonstrating its orthogonal applicability. We applied the CSD ($S,S$) loss to ImitKD (Lin et al., 2020), GKD (Agarwal et al., 2024), and DistiLLM (Ko et al., 2024). The primary distinction among these methods, apart from their losses, lies in the choice of dataset $\mathcal{D}$: ImitKD uses purely student-generated on-policy data, GKD combines fixed data with student outputs, and DistiLLM adaptively selects between them based on validation loss. As a result, the average ROUGE-L score improved for both GPT-2-0.1B and GPT-2-0.3B students across all cases. The best result on each benchmark was also achieved by our method, with particularly strong performance under pure on-policy settings. We also evaluated using GPT-4 as the judge in Figure˜4, where our best model was judged superior to other baselines. There may exist CSD variants other than ($S,S$) that perform better for specific $\mathcal{D}$, but we leave this exploration to future work. Finally, applying our best setting to a larger OpenLLaMA also outperformed recent baselines, demonstrating the scalability of CSD with respect to model size. We also provide comparisons with more baselines in Table˜5 of Appendix˜D.

### v2

# arXiv:2509.25837v2[cs.LG] 01 Mar 2026

@@ Abstract
Large language models (LLMs) deliver remarkable performance but are costly to deploy, motivating knowledge distillation (KD) for efficient inference. Existing KD objectives typically match student and teacher probabilities via softmax, which blurs valuable logit information. While direct logit distillation (DLD) mitigates softmax smoothing, it fails to account for logit shift invariance, thereby restricting the solution space. We propose Concrete Score Distillation (CSD), a discrete score-matching objective that overcomes both softmax-induced smoothing and restrictions on the optimal solution set. We resolve the training instability and quadratic complexity of discrete score-matching in autoregressive LLMs, and the resulting CSD objective aligns relative logit differences across all vocabulary pairs between student and teacher with flexible weighting. We provide both mode-seeking and mode-covering instances within our framework and evaluate CSD on task-agnostic instruction-following, task-specific, and general chat capability distillation using GPT-2-1.5B, OpenLLaMA-7B, Gemma-7B-IT, Qwen2.5-7B-IT, and Gemma2-9B-IT teachers. Experiments show that CSD consistently surpasses recent KD objectives, achieves favorable fidelity–diversity trade-offs, demonstrating its scalability and effectiveness for LLM distillation. Code: https://github.com/aailab-kaist/CSD.
@@ 1 Introduction
In traditional KD, direct logit distillation (DLD) (Ba and Caruana, 2014; Urban et al., 2017) has been proposed as an alternative strategy, with advantages in generalization capability and in removing the softmax smoothing (Kim et al., 2021). However, such approaches have not been thoroughly explored in the context of LLMs. This paper identifies a key drawback of DLD: its restriction on the optimal solution set as described in Figure˜1(c). Considering the softmax activation in inference, it is sufficient for the teacher’s and student’s logits to agree up to an additive constant, but the previous solutions of DLD fail to accommodate such an acceptable slack constant, a.k.a. logit shift invariance. Such a restriction on the solution set may hinder the discovery of optimal solutions in distillation, particularly when the teacher and student models have a large capacity gap, as is often the case with LLMs. Therefore, the goal of this paper is to establish a design space of distillation losses that overcome both the softmax-induced smoothing of teacher knowledge and the restriction on the solution set.
This paper adopts the idea from energy-based models (Song and Kingma, 2021), which design objectives that avoid the constraint of probabilistic models (sum-to-one) by using the score-matching objective (Hyvärinen and Dayan, 2005). We propose Concrete Score Distillation (CSD), a discrete form of the score-matching objective (Meng et al., 2022) adapted for autoregressive LLM distillation. We address training instability and computational overhead arising when applying the score-matching objective to LLMs, and provide theoretical guarantees of optimality, showing that its solution set is broader than that of DLD from both theoretical and empirical perspectives. The resulting objective reduces to matching the relative logit differences across all pairs of vocabulary items between the student and teacher, while allowing flexible weighting across all vocabulary pairs in linear time with respect to vocabulary size. Furthermore, we present instances within our framework that exhibit both mode-seeking and mode-covering properties.
In our experiments, we conducted task-agnostic instruction-following distillation, task-specific distillations (summarization, mathematics, and translation), and general chat capability distillation using GPT-2 (Radford et al., 2019), OpenLLaMA (Geng and Liu, 2023), Gemma (Team et al., 2024), Qwen2.5 (Team and others, 2024), and Gemma2 (Team et al., 2024) backbones. The proposed CSD consistently outperformed recent probability-matching objectives as well as direct logit distillation. By appropriately choosing weighting functions, we further demonstrated that our method resides on the frontier of the diversity–fidelity trade-off. Finally, we observed complementary performance gains when integrating our loss with on-policy techniques.
@@ 2.1 Knowledge Distillation of Large Language models
We consider autoregressive large language models (LLMs), consisting of a teacher $p_{T}$ and a student $q_{\theta}$ with $\theta\in\Theta$, where the student is a smaller and more efficient model. Given an input context $\mathbf{c}$}, the student generates an output sequence ${\mathbf{y}}=(y_{1},y_{2},\dots,y_{L})$ with probability $q_{\theta}({\mathbf{y}}|{\mathbf{c}})=\prod_{t=1}^{L}{q_{\theta}(y_{t}|{\mathbf{c}},{\mathbf{y}}_{<t})}$, where $L$ denotes the sequence length, and the teacher’s probability is defined analogously. Each token $y_{t}$ is drawn from the fixed vocabulary set $\mathcal{V}:=\{v_{1},v_{2},$...\}. As in prior works (Lin et al., 2020; Ko et al., 2024), we assume the teacher and student share the same vocabulary set. To compute the token probability $q_{\theta}(y_{t}|{\mathbf{c}},{\mathbf{y}}_{<t})$, an LLM typically adopts a parametric function $f_{\theta}:{\mathcal{V}}^{|{\mathbf{c}}|}\times{\mathcal{V}}^{t-1}\rightarrow\mathbb{R}^{|{\mathcal{V}}|}$, which maps the input $({\mathbf{c}},{\mathbf{y}}_{<t})$ to a logit vector $f_{\theta}({\mathbf{c}},{\mathbf{y}}_{<t})\in\mathbb{R}^{|{\mathcal{V}}|}$. The logit corresponding to token $y_{t}$ is denoted by $f_{\theta}({\mathbf{c}},{\mathbf{y}}_{<t})[y_{t}]$. For brevity of notation, the input arguments of the function $f_{\theta}$ will be omitted hereafter. Let $f_{T}$ be the parametric function of the teacher. Accordingly, the probability of each token is calculated through the softmax transformation:
@@
However, $D_{\text{KL}}$ focuses on the teacher’s probabilities and is constrained by the softmax. As shown in Figure˜1(b), although the teacher carries rich knowledge across all vocabulary items at the logit level, much of it is lost after softmax, and the teacher provides nearly identical gradient signals to most minor tokens. Accordingly, in classical KD studies (Ba and Caruana, 2014; Urban et al., 2017), direct logit distillation (DLD) has been widely adopted as a logit-level mean squared error (MSE) loss:
@@ 2.2 Score Matching for a discrete random variable
Score-matching (SM) (Hyvärinen and Dayan, 2005) was originally proposed in energy-based models (Song and Kingma, 2021) with continuous variables ${\mathbf{x}}\in\mathbb{R}^{d}$. An energy function $E_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ maps $\mathbf{x}$} to a scalar. The corresponding probability and the score-matching objective are given by:
@@
where $p_{\text{data}}$ is the data distribution, $Z_{\theta}=\int_{{\mathbf{x}}}\text{exp}(-E_{\theta}({\mathbf{x}}))\text{d}{\mathbf{x}}$ is the partition function, and $w(\cdot)$ is a weighting function. The term $\nabla_{{\mathbf{x}}}\log{q_{\theta}({\mathbf{x}})}=-\nabla_{{\mathbf{x}}}E_{\theta}({\mathbf{x}})$ is known as the Stein score, which uniquely identifies the probability distribution without requiring the computation of $Z_{\theta}$. $\mathcal{L}_{\text{SM}}$ facilitates the design of losses without considering the normalization constraint of probabilistic models. The probability computation $q_{\theta}$ here follows, analogously, the form of the LLM probabilities in Eq.˜1. The difference is that an LLM outputs energy values $f_{\theta}$ over all finite states at once, whereas an EBM handles continuous variables, so that each input to $E_{\theta}$ yields only a single scalar output.
@@
where $p_{\text{data}}$ is the data distribution defined over a discrete state, and $w(\cdot,\cdot)$ is a positive weighting function. Previous work on language models (Lou et al., 2024) typically adopted this loss by directly parameterizing the concrete score (also known as discrete diffusion models) to mimic the data distribution. In contrast, we take this concept as a starting point to design logit-level distillation losses for autoregressive-type language models, which are more dominant in real-world applications.
@@ 3.1 Concrete Score Distillation for Large Language Models
Training instability is a well-known issue in likelihood ratio estimation (Rhodes et al., 2020). Following Higuchi and Suzuki (2025), we address it by applying a monotonically increasing function to the concrete scores. In particular, we adopt the logarithm, which yields the following objective:
@@
(Consistency) Given context $\mathbf{c}$} and prefix ${\mathbf{y}}_{<t}$, assume model capacity $|\Theta|\rightarrow\infty$. For any $w(\cdot,\cdot)>0$, define the set of optimal parameters as $\Theta_{\text{CSD}}^{*}=\operatorname*{arg\,min}_{\theta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\in\Theta}}{\mathcal{L}_{\text{CSD}}\left(\theta;p_{T},w\right)}$. Then, for any $\theta^{*}\in\Theta_{\text{CSD}}^{*}$, we have $\mathcal{L}_{\text{CSD}}\left(\theta^{*};p_{T},w\right)=0$, and the following holds for all $y_{t}\in\mathcal{V}$:
@@
(Solution Superset) Assume model capacity $|\Theta|\rightarrow\infty$, let the set of optimal parameters $\Theta_{\text{CSD}}^{*}=\operatorname*{arg\,min}_{\theta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\in\Theta}}{\mathcal{L}_{\text{CSD}}\left(\theta;p_{T},w\right)}$ and $\Theta_{\text{DLD}}^{*}=\operatorname*{arg\,min}_{\theta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\in\Theta}}{\mathcal{L}_{\text{DLD}}\left(\theta;p_{T},w\right)}$, then following holds:
@@ 3.2 Gradient Computation and Analysis
1 Compute the student logit $f_{\theta}[y_{t}]=f_{\theta}({\mathbf{c}},{\mathbf{y}}_{<t})[y_{t}],\forall y_{t}\in\mathcal{V}$.
@@
3  Compute the teacher logit $f_{T}[y_{t}]=f_{T}({\mathbf{c}},{\mathbf{y}}_{<t})[y_{t}],\forall y_{t}\in\mathcal{V}$.
@@ 7   Compute the weighted normalized logits:
The proof is provided in Section˜A.3. For the actual training time and memory usage, please refer to Table˜9 in Appendix˜D. These results follow from factorizing the independent variables. Algorithm˜1 further details the gradient computation of Eq.˜9 step by step, with each step requiring only linear time over the vocabulary. An alternative approach is to use Monte Carlo estimation as described in Algorithm˜2 of Appendix˜C. Instead of taking a weighted sum over all possible states of $y_{t}$ with $w_{1}$, one can draw a single sample of $y_{t}$ according to probability $w_{1}$ and compute the loss in expectation. The Monte Carlo estimation, unlike the analytic gradient form, does not require assuming independence between the two variables of $w$, allowing it to model the joint weighting function space directly. However, defining a joint weighting function over two discrete vocabulary spaces is generally difficult. We found that independent weighting functions capture various behaviors in Section˜4.4. In these cases, Monte Carlo estimation increases the variance within batched samples, which slightly slows the convergence compared to the analytic computation (see Figure˜5(c)).
@@ 4 Experiments
This section comprehensively validates the effectiveness of the proposed Concrete Score Distillation (CSD) across various experimental setups. Section˜4.1 shows results on task-agnostic instruction-following distillation, comparing CSD with alternative loss functions and assessing its performance when combined with on-policy methods. Section˜4.2 further examines task-specific settings, including math, summarization, and translation, to evaluate the applicability of CSD. Section˜4.3 evaluates scalability by examining whether general conversational abilities can also be distilled. Finally, Section˜4.4 establishes the contribution of each component in CSD through ablation studies.
@@ 4.1 Task-Agnostic Instruction-Following Distillation
We follow the training setup of DistiLLM (Ko et al., 2024). For the distillation dataset $\mathcal{D}$, we use databricks-dolly-15k (Conover et al., 2023). We first fine-tune the teacher on this dataset and then distill it into student models. We use the detached student probability as the default choice for both $w_{1}$ and $w_{2}$, and we apply it similarly to the weights in DLD. Please refer to Section˜C.1 for further details on backbone, training configuration, baseline, and the evaluation protocol.
@@
(b) Decoding temperature ablation in $[0.2,1.8]$.
@@
Table 2: Instruction-following performance of CSD with on-policy techniques for various backbones. $\mathcal{D}$ denotes the distillation dataset. ROUGE-L scores are averaged over five random seeds, with the best score for each student highlighted in bold. CSD and DLD use the student probability weighting.
@@
Table 3: Task-specific distillation from Gemma-7B-IT to Gemma-2B-IT.
Summarization Translation GSM8K Loss ROUGE-L COMET Accuracy Teacher 37.09 79.23 60.27 KL 35.02 73.96 24.03 JS 35.60 74.05 23.73 TV 27.49 73.73 0.00 Jeffrey 35.29 74.02 23.28 SKL 25.86 59.65 0.00 SRKL 26.68 73.10 0.00 RKL 0.00 45.02 0.00 DLD $(S)$ 0.00 21.52 0.00 DLD-max $(T)$ 32.54 65.28 17.74 CSD $(T,S)$ 35.67 74.14 25.78
Loss-level comparison: To purely analyze the effect of the distillation loss itself, this comparison excludes the use of pretraining losses, initialization with an SFT-tuned student, and any on-policy techniques. Table˜1 shows that the proposed CSD objective outperforms the other nine objectives, ranking first on three of the five benchmarks, second on one, and achieving the highest average score. SKL (Ko et al., 2024) and AB (Wang et al., 2025) exhibit slightly lower performance than previously reported, likely due to their reliance on pretraining losses or on-policy techniques. Figure˜3(a) shows the fidelity–diversity trade-off based on ROUGE-L and Self-BLEU scores. Traditionally, KL favors diversity, whereas RKL favors mode-seeking. Within this trade-off, SKL, SRKL, TV, and AB achieve higher ROUGE-L scores than RKL, but at the cost of reduced diversity, reflecting a stronger emphasis on fidelity. Diversity, however, remains an important aspect of user experience in instruction-following, and it becomes a valuable metric as it enhances performance when combined with best-of-N sampling. The proposed CSD provides an additional lever to control the fidelity–diversity trade-off. By default, using the detached student probabilities ($S,S$) yields the highest fidelity. Replacing one side with uniform ($U,S$) or with the teacher ($T,S$) gradually increases diversity. This is likely because the ($S,S$) makes the model focus only on regions where the student already assigns a high likelihood, limiting its exploratory ability. The trade-off offered by CSD envelopes those of existing losses, and we expect that even better operating points may exist within the design space of $w_{1}$ and $w_{2}$. Figure˜3(b) presents an ablation on temperature, which enables easy adjustment of the trade-off during inference. Even within a reasonable range of decoding temperature, CSD achieves better trade-off points than other losses. In particular, CSD ($U,S$) demonstrates a well-balanced exchange between diversity and fidelity through temperature adjustment.
Orthogonal improvement with recent on-policy advances: Table˜2 reports the performance of recent distillation baselines augmented with the CSD loss, demonstrating its orthogonal applicability. We applied the CSD ($S,S$) and DLD ($S$) loss to ImitKD (Lin et al., 2020), GKD (Agarwal et al., 2024), and DistiLLM (Ko et al., 2024). DLD-mean refers to the mean-centered DLD variant, as described in Section˜A.4. The primary distinction among these methods, apart from their losses, lies in the choice of dataset $\mathcal{D}$: ImitKD uses purely student-generated on-policy data, GKD combines fixed data with student outputs, and DistiLLM adaptively selects between them based on validation loss. As a result, the average ROUGE-L score improved for both GPT-2-0.1B and GPT-2-0.3B students in all settings, compared to both the baseline and the corresponding DLD versions. The best result on each benchmark was also achieved by our method, with particularly strong performance under pure on-policy settings. We also evaluated using GPT-4 as the judge in Figure˜4, where our best model was judged superior to other baselines. There may exist CSD variants other than ($S,S$) that perform better for specific $\mathcal{D}$, but we leave this exploration to future work. Finally, applying our best setting to a larger OpenLLaMA also outperformed baselines, demonstrating the scalability of CSD with respect to model size. We provide comparisons with more baselines in Table˜7 of Appendix˜D.